# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Accident Prediction

### v3

@@ Abstract
Traffic accidents present substantial challenges to human safety and socio-economic development in urban areas. Developing a reliable and responsible traffic accident prediction model is crucial to addressing growing public safety concerns and enhancing the safety of urban mobility systems. Traditional methods face limitations at fine spatiotemporal scales due to the sporadic nature of high-risk accidents and the predominance of non-accident characteristics. Furthermore, while most current models show promising occurrence prediction, they overlook the uncertainties arising from the inherent nature of accidents, and then fail to adequately map the hierarchical ranking of accident risk values for more precise insights. To address these issues, we introduce the Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNN), the first uncertainty-aware probabilistic graph deep learning model in road-level traffic accident prediction for multi-steps. This model integrates the interpretability of the statistical Tweedie family model and the expressive power of graph neural networks. Its decoder innovatively employs a compound Tweedie model—a Poisson distribution to model the frequency of accident occurrences and a Gamma distribution to assess injury severity, supplemented by a zero-inflated component to effectively identify exessive non-incident instances. Empirical tests using real-world traffic data from London, UK, demonstrate that the STZITD-GNN surpasses other baseline models across multiple benchmarks and metrics, including accident risk value prediction, uncertainty minimisation, non-accident road identification and accident occurrence accuracy. Our study demonstrates that STZTID-GNN can effectively inform targeted road monitoring, thereby improving urban road safety strategies.
@@
Spatiotemporal Sparse Data Mining , Traffic Accident Prediction , Uncertainty Quantification , Zero-Inflated Tweedie Distribution
@@ 1 Introduction
Traffic accidents persist as a major barrier to sustainable urban development (Dai et al.,, 2018; Tu et al.,, 2020; Barahimi et al.,, 2021; Rodrigue,, 2020). The recent World Health Organisation (WHO) report reveals that road traffic accidents claimed approximately 1.19 million lives worldwide in 2021, making them the leading cause of mortality and disability and imposing an economic burden of up to US\$1.8 trillion, approximately 10–12% of the global Gross Domestic Product (GDP)(Organization,, 2023).
@@
To support reductions in traffic accidents, there is a growing emphasis on the development of models aimed at predicting the location, severity, and causes of accidents. The early approaches relied mainly on linear regression models (Mountain et al.,, 1996; Greibe,, 2003), models that incorporate random parameters (El-Basyouny and Sayed,, 2009; Dinu and Veeraragavan,, 2011), and geographically weighted regression (GWR) models (Li et al.,, 2013; Xu and Huang,, 2015) to predict regional traffic accidents. Recent research has turned to deep learning-based methods, notably graph neural networks (GNNs), which are able to handle large datasets and capture complex spatial and temporal representations of urban road networks. Examples include time-varying GNN (Zhou et al., 2020b, ), multiview fusion GNN(Trirat et al.,, 2023) and subgraph-based GNN(Zhao et al.,, 2023). Despite substantial advancements in model development, accurately predicting road accidents at a fine spatiotemporal granularity road presents several challenges:
@@
  • Spatially Imbalanced Accident Occurrence:The prediction of road accidents is challenging due to the uneven spatial distribution of accidents. Firstly, zero-inflation is inherent to accident data, where areas with no accidents are much more common than those with accidents. This spatially skewed distribution causes models to be biased toward predicting zero accidents, complicating the effective training of predictive algorithms(Wu et al.,, 2022). Secondly, where accidents do occur, they are predominantly minor, leading to an abundance of low-risk scores across the spatial grid. This creates a misleading representation that underestimates the potential severity of less frequent, more severe accidents(Shirazi and Lord,, 2019; Saha et al.,, 2020). The combination of these factors—widespread zero-accident areas and the dominance of minor accidents in accident-prone zones—results in a challenging scenario for models that need to understand and predict the full spectrum of accident severities across different areas (Wang et al.,, 2023).
@@
  • Temporal Constraints with Uncertainty: Building on the spatial challenges, the temporal dynamics of traffic accidents introduce additional complexity. Particularly, predictive models for traffic accidents face the challenge of accurately forecasting not just immediate but also future conditions, necessitating reliable multi-step predictions (Soltani and Qadikolaei,, 2024). The nature of traffic accidents, characterized by their non-linearity and the variable time intervals between events, adds complexity to this task. As the prediction horizon extends, uncertainty increases, particularly because future traffic conditions and potential accident triggers can change substantially over time(Qian et al.,, 2022). Most current deep learning methods do not fully address these challenges. They tend to overlook the complexities of aleatoric uncertainty, which arises from the inherent randomness in the data. These models often rely on assumptions of uniform variance across all data points, producing point estimates from average outcomes(Wang et al.,, 2023; Zhuang et al.,, 2022; Jiang et al.,, 2023). This approach fails to capture the temporal variability and the unpredictable nature of accidents, resulting in predictions that do not adequately reflect the true range of future possibilities.
@@
To address the intricate challenges posed by the spatial and temporal distribution of accident data, we introduce the Spatio-temporal Zero-Inflated Tweedie Graph Neural Network (STZITD-GNN). This model combines spatio-temporal graph machine learning with statistical methods, enabling enhanced uncertainty quantification over multi-step prediction intervals. It is specifically designed to handle the probabilistic assessment of non-recurrent, road-level traffic accidents, providing a robust solution to predict spatially and temporally varying accident data.
@@
The proposed framework integrates a cohesive encoder that combines a Gated Recurrent Unit (GRU) for capturing temporal dynamics and a Graph Attention Network (GAT) for spatial relationships, ensuring a synchronised capture of spatio-temporal dependencies across individual roads and time steps. The decoder leverages a Tweedie (TD)-based distribution, a flexible compound Poisson-gamma model that simultaneously models both exact zero occurrences and continuous positive values. The choice of the TD distribution is motivated by its established efficacy in fields challenged by nonnormal distributions, such as accident frequency estimation (Wang et al.,, 2019), insurance claims (Smyth and Jørgensen,, 2002), actuarial studies (Shi,, 2016), and meteorological precipitation (Dunn,, 2004). This approach allows the model to output the TD distribution parameters, explicitly determining the upper and lower bounds for the prediction intervals of each sample, thus providing a probabilistic representation of each road’s accident risk scores.
@@
Given the inherent extreme imbalance that traditional TD models struggle to adequately capture (Khosravi et al.,, 2010; Abe and Yadohisa,, 2017), we have enhanced our model by incorporating an additional parameter, $\pi$, to estimate the probability mass of zero values. This adjustment leads to the creation of the Zero-Inflated Tweedie (ZITD) distribution on the decoder, which offers a more comprehensive and responsible depiction of real accident scenarios. This refinement ensures that our model not only captures the complexities of traffic accident data, but also enhances the accuracy and reliability of its predictions. Specifically, our main contributions can be summarised as follows.
@@
  • An innovative ZITD distribution is employed within our model to effectively address the uncertainties associated with zero-inflation and the long-tailed characteristics prevalent in road accident data. This approach enhances the model’s capability to accurately represent the complex distribution patterns observed in traffic accident data.
  • To the best of our knowledge, the STZITD-GNN is the first work to use probabilistic framework for granular, multistep road-level accident forecasting. The use of Likelihood-based Loss minimization rather than the tradional regression ones enhances its uncertainty-aware predictive accuracy and reliability.
@@ 2 Literature Review
Predicting traffic accidents is essential for urban planning and mobility management. Initially, research in this area predominantly utilised traditional statistical methods and classical machine learning techniques, such as regression models (Chang,, 2005; Lord et al.,, 2005; Caliendo et al.,, 2007; Bergel-Hayat et al.,, 2013), Bayesian networks (Martin et al.,, 2009; Mujalli and De Oña,, 2011; Hossain and Muromachi,, 2012), tree-based models (Chong et al.,, 2004; Chang and Chen,, 2005; Wang et al.,, 2010; Lin et al.,, 2015), and the k-nearest-neighbour method (Sayed and Abdelwahab,, 1998; Tang and Gao,, 2005; Lv et al.,, 2009). These early studies were generally confined to small geographical areas and demonstrated limited effectiveness in capturing the nonlinear and dynamic spatio-temporal patterns of the contributing factors (Chang and Chen,, 2005; Anderson,, 2009; Zhang et al.,, 2014). In addition, these methods often analyse accident data in isolation, neglecting critical interdependencies between different locations (Wang et al.,, 2013; Wang et al., 2021a, ). As a result, their broader applicability to city-wide analysis with large data size was constrained, highlighting the need for more advanced models capable of addressing these limitations.
@@
Recent advances in deep learning techniques have shown promise in urban traffic accident prediction by incorporating spatio-temporal characteristics of the data. Convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph-based networks, such as graph-convolution networks (GCNs), have demonstrated their effectiveness in capturing intricate spatial patterns, capturing temporal dependencies, and modelling complex interactions (Wang and Cao,, 2021; Liu et al.,, 2020). As summarised in Table 1, Chen et al., (2016) were among the first to explore city level accident prediction employing a stack denoise autoencoder. They integrated human mobility GPS data and historical incident point data at the grid level to map the real-time accident situation in Tokyo. However, while their approach attempted to capture a larger spatial area, it did not consider urban geo-semantic information for precise and long-term accident prediction. Later, Chen et al., (2018) introduced a Stack Denoise Convolutional Autoencoder that incorporated spatial dependencies using stacked CNNs. Nonetheless, both of these studies overlooked the influence of temporal factors. Although RNNs are capable of handling temporal information, they are more suitable for short-term temporal learning (Ren et al.,, 2018). To better address the temporal aspect, Ren et al., (2018) used Long-Short-Term Memory (LSTM) to consider influential temporal factors between several locations. Their work represented the first attempt to incorporate temporal factors into deep learning models for the prediction of city-wide traffic accidents. Similarly, Moosavi et al., (2019) employed LSTM networks for country-wide accident prediction, demonstrating their efficacy in predicting a binary outcome. Building on the LSTM framework, Yuan et al., (2018) proposed a more advanced model known as the Convolutional Long Short-Term Memory (Hetero-ConvLSTM) neural network based on spatial heterogeneous data. This model aimed to jointly address spatial heterogeneity and capture temporal features in state-wide accident prediction. However, their approach relied on pre-defined moving windows based on the spatial features of Iowa, United States, limiting its generalisability to other cities. Bao et al., (2019) integrated CNNs, LSTM as well as ConvLSTM into a spatiotemporal convolutional long short-term memory network (STCL-Net). Their work presented a more practical and feasible approach for the prediction of city-wide traffic accidents. By combining the strengths of these models, STCL-Net effectively captured spatiotemporal dependencies in accident data and influential factors, leading to improved prediction performance.
@@
To capture local and global dynamics with hierarchical spatial information, Zhu et al., (2019) proposed the Deep Spatio-temporal Attention Learning Framework, which incorporates a spatio-temporal attention mechanism to effectively capture the dynamic impact of traffic accidents at different spatial levels. One notable aspect of their work was the emphasis on using real traffic administrative areas instead of manually divided grids. However, this approach does not incorporate spatial relationships, which limits its capacity to model the spatial heterogeneities that characterise traffic accidents. In a similar vein, Huang et al., (2019) introduced the Deep Dynamic Fusion Network (DFN) framework, which aggregates heterogeneous external factors across both spatial and temporal dimensions. Using a temporal aggregation layer, the DFN framework is designed to automatically capture external influences from latent temporal dimensions.
@@
Although the mentioned studies effectively captured temporal and spatial patterns in traffic accident data, yielding promising prediction outcomes, their lack of consideration of latent spatial correlations within urban structures can disrupt regional patterns in practical urban science and lead to prediction inaccuracies. Therefore, Graph-based networks have become a key method for city-wide traffic accident prediction by leveraging spatial graph features from actual urban networks (Xue et al.,, 2022; Cheng et al.,, 2022; Li et al., 2022a, ; Li et al., 2022b, ).
@@
Zhou et al., 2020b developed the RiskOracle framework by constructing an urban graph of regions and integrating the Differential Time-varying Graph Neural Network (DTGN) to identify dynamic correlations between historical traffic conditions and risks, enabling simultaneous predictions of traffic flow and accidents. To tackle zero-inflation in accident data, they introduced a data enhancement strategy based on a priori knowledge (PKDE), defining subregional accident labels for training. This approach adjusts zero accident values to a range between 0 and 1, maintaining the hierarchy of actual accident levels. Expanding on this foundation, Zhou et al., 2022b refined the DTGN framework by embedding an LSTM-based hierarchical sequence learning architecture, accommodating multiscale historical spatio-temporal data for nuanced long-term and short-term predictions. Wang et al., 2021a and Wang et al., 2021b both focus on the integration of geographical and semantic relationships through different graph structures. While Wang et al., 2021a used accident, road, and POI graphs to capture regional semantics and introduced modules for geographical and semantic spatio-temporal correlations, Wang et al., 2021b ) combined attention-based LSTM, CNN, and GCN for multi-view predictions of traffic accidents at various scales. Although both approaches tackled the zero-inflation issue with promising results, they share limitations in addressing dynamic geographical semantics and alignment issues in graph structures. In contrast, Trirat et al., (2023) proposed a multi-view graph neural network incorporating dynamic and static similarity information, offering a more dynamic approach to understanding traffic accidents. Using the Huber loss function, a parameter-based regression loss, their model aimed at robust regression that effectively accommodates zero inflation, showcasing an advanced strategy for precise prediction of urban traffic accidents.
@@
Recent attempts to apply graph-based deep learning for road-level accident prediction have shown promise, but also highlight areas for improvement. Yu et al., (2021) developed a Deep Spatiotemporal Graph Convolutional Network with a unique three-layer network, treating the road graph, spatio-temporal data, and embeddings separately. They tackled zero inflation through undersampling, balancing the dataset between risky and non-risky roads, and framing accident prediction as a binary classification. While this method effectively predicts accidents at the road level, the binary outcome overlooks the varying degrees of accident risk across roads. Similarly, Wu et al., (2022) designed a fully connected dynamic network to predict road-level accidents by focussing on dynamic spatio-temporal correlations. To mitigate zero inflation, they introduced a cost-sensitive learning module that enhances positive sample classification accuracy by considering the imbalance between zero and non-zero accident values. However, this work only considered limited road segments in New York City with binary accident occurrence outputs, highlighting the need for further development to capture the full spectrum of accident severity and provide a more detailed understanding of road-level accident scenarios.
@@
In summary, despite significant progress in urban traffic accident prediction yielding promising results, existing studies have yet to fully address uncertainty quantification in their predictions and provide multi-step accident assessments at the detailed road level.
@@
Table 1: Summary of Deep Learning Methods in Traffic Accident Prediction
@@
Chen et al., (2016) Stack Denoise Autoencoder Regions N/A Single Accident Risk Score
Ren et al., (2018) Long Short-term Memory network Regions N/A Multiple Accident Risk Score
Chen et al., (2018) Stack Denoise Convolutional Autoencoder Regions N/A Single Accident Risk Score
Yuan et al., (2018) Convolutional Long Short-Term Memory Regions N/A Multiple Accident Risk Score
Moosavi et al., (2019) Long Short-term Memory Network Regions N/A Single Accident Occurrence
Zhu et al., (2019) Deep Spatial-Temporal Attention Learning Framework Regions N/A Multiple Accident Risk Score
Huang et al., (2019) Dynamic Fusion Network Regions N/A Single Accident Occurrence
Bao et al., (2019) Spatiotemporal Convolutional LSTM Regions N/A Single Accident Risk Score
Zhou et al., 2020b DTGN with PKDE Regions Yes Single Accident Risk Score
Wang et al., 2021a GSNet with Weighted Loss Regions Yes Single Accident Risk Score
Wang et al., 2021b MVMT-ST Networks Regions Yes Single Accident Risk Score
Yu et al., (2021) Deep ST-GCN with Undersampling Roads Yes Single Accident Occurrence
Wu et al., (2022) MADGCN with Cost-sensitive Loss Roads Yes Multiple Accident Occurrence
Zhou et al., 2022b Enhanced DTGN Regions Yes Multiple Accident Risk Score
Trirat et al., (2023) MV-GNNs with Huber Loss Regions Yes Multiple Accident Risk Score

### v4

# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Crash Prediction

@@ Abstract
Traffic crashes present substantial challenges to human safety and socio-economic development in urban areas. Developing a reliable and responsible traffic crash prediction model is crucial to addressing growing public safety concerns and enhancing the safety of urban mobility systems. Traditional methods face limitations at fine spatiotemporal scales due to the sporadic nature of high-risk crashes and the predominance of non-crash characteristics. Furthermore, while most current models show promising occurrence prediction, they overlook the uncertainties arising from the inherent nature of crashes, and then fail to adequately map the hierarchical ranking of crash risk values for more precise insights. To address these issues, we introduce the Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNN), the first uncertainty-aware probabilistic graph deep learning model in road-level traffic crash prediction for multi-steps. This model integrates the interpretability of the statistical Tweedie family model and the expressive power of graph neural networks. Its decoder innovatively employs a compound Tweedie model—a Poisson distribution to model the frequency of crash occurrences and a Gamma distribution to assess injury severity, supplemented by a zero-inflated component to effectively identify exessive non-incident instances. Empirical tests using real-world traffic data from London, UK, demonstrate that the STZITD-GNN surpasses other baseline models across multiple benchmarks and metrics, including crash risk value prediction, uncertainty minimisation, non-crash road identification, and crash occurrence accuracy. Our study demonstrates that STZTID-GNN can effectively inform targeted road monitoring, thereby improving urban road safety strategies.
@@
Spatiotemporal Sparse Data Mining , Traffic Crash Prediction , Uncertainty Quantification , Zero-Inflated Tweedie Distribution
@@ 1 Introduction
Traffic crashes persist as a major barrier to sustainable urban development (Dai et al.,, 2018; Tu et al.,, 2020; Barahimi et al.,, 2021; Rodrigue,, 2020). The recent World Health Organisation (WHO) report reveals that road traffic crashes claimed approximately 1.19 million lives worldwide in 2021, making them the leading cause of mortality and disability and imposing an economic burden of up to US\$1.8 trillion, approximately 10–12% of the global Gross Domestic Product (GDP)(Organization,, 2023).
@@
To support reductions in traffic crashes, there is a growing emphasis on the development of models aimed at predicting the location, severity, and causes of crashes. The early approaches relied mainly on linear regression models (Mountain et al.,, 1996; Greibe,, 2003), models that incorporate random parameters (El-Basyouny and Sayed,, 2009; Dinu and Veeraragavan,, 2011), and geographically weighted regression (GWR) models (Li et al.,, 2013; Xu and Huang,, 2015) to predict regional traffic crashes. Recent research has turned to deep learning-based methods, notably graph neural networks (GNNs), which are able to handle large datasets and capture complex spatial and temporal representations of urban road networks. Examples include time-varying GNN (Zhou et al., 2020b, ), multiview fusion GNN(Trirat et al.,, 2023) and subgraph-based GNN(Zhao et al.,, 2023). Despite substantial advancements in model development, accurately predicting road crashes at a fine spatiotemporal granularity road presents several challenges:
@@
  • Spatially Imbalanced Crash Occurrence:The prediction of road crashes is challenging due to the uneven spatial distribution of crashes. Firstly, zero-inflation is inherent to crash data, where areas with no crashes are much more common than those with crashes. This spatially skewed distribution causes models to be biased toward predicting zero crashes, complicating the effective training of predictive algorithms(Wu et al.,, 2022). Secondly, where crashes do occur, they are predominantly minor, leading to an abundance of low-risk scores across the spatial grid. This creates a misleading representation that underestimates the potential severity of less frequent, more severe crashes(Shirazi and Lord,, 2019; Saha et al.,, 2020). The combination of these factors—widespread zero-crash areas and the dominance of minor crashes in crash-prone zones—results in a challenging scenario for models that need to understand and predict the full spectrum of crash severities across different areas (Wang et al.,, 2023).
@@
  • Temporal Constraints with Uncertainty: Building on the spatial challenges, the temporal dynamics of traffic crashes introduce additional complexity. Particularly, predictive models for traffic crashes face the challenge of accurately forecasting not just immediate but also future conditions, necessitating reliable multi-step predictions (Soltani and Qadikolaei,, 2024). The nature of traffic crashes, characterized by their non-linearity and the variable time intervals between events, adds complexity to this task. As the prediction horizon extends, uncertainty increases, particularly because future traffic conditions and potential crash triggers can change substantially over time(Qian et al.,, 2022). Most current deep learning methods do not fully address these challenges. They tend to overlook the complexities of aleatoric uncertainty, which arises from the inherent randomness in the data. These models often rely on assumptions of uniform variance across all data points, producing point estimates from average outcomes(Wang et al.,, 2023; Zhuang et al.,, 2022; Jiang et al.,, 2023). This approach fails to capture the temporal variability and the unpredictable nature of crashes, resulting in predictions that do not adequately reflect the true range of future possibilities.
@@
To address the intricate challenges posed by the spatial and temporal distribution of crash data, we introduce the Spatio-temporal Zero-Inflated Tweedie Graph Neural Network (STZITD-GNN). This model combines spatio-temporal graph machine learning with statistical methods, enabling enhanced uncertainty quantification over multi-step prediction intervals. It is specifically designed to handle the probabilistic assessment of non-recurrent, road-level traffic crashes, providing a robust solution to predict spatially and temporally varying crash data.
@@
The proposed framework integrates a cohesive encoder that combines a Gated Recurrent Unit (GRU) for capturing temporal dynamics and a Graph Attention Network (GAT) for spatial relationships, ensuring a synchronised capture of spatio-temporal dependencies across individual roads and time steps. The decoder leverages a Tweedie (TD)-based distribution, a flexible compound Poisson-gamma model that simultaneously models both exact zero occurrences and continuous positive values. The choice of the TD distribution is motivated by its established efficacy in fields challenged by nonnormal distributions, such as crash frequency estimation (Wang et al.,, 2019), insurance claims (Smyth and Jørgensen,, 2002), actuarial studies (Shi,, 2016), and meteorological precipitation (Dunn,, 2004). This approach allows the model to output the TD distribution parameters, explicitly determining the upper and lower bounds for the prediction intervals of each sample, thus providing a probabilistic representation of each road’s crash risk scores.
@@
Given the inherent extreme imbalance that traditional TD models struggle to adequately capture (Khosravi et al.,, 2010; Abe and Yadohisa,, 2017), we have enhanced our model by incorporating an additional parameter, $\pi$, to estimate the probability mass of zero values. This adjustment leads to the creation of the Zero-Inflated Tweedie (ZITD) distribution on the decoder, which offers a more comprehensive and responsible depiction of real crash scenarios. This refinement ensures that our model not only captures the complexities of traffic crash data, but also enhances the accuracy and reliability of its predictions. Specifically, our main contributions can be summarised as follows.
@@
  • An innovative ZITD distribution is employed within our model to effectively address the uncertainties associated with zero-inflation and the long-tailed characteristics prevalent in road crash data. This approach enhances the model’s capability to accurately represent the complex distribution patterns observed in traffic crash data.
  • To the best of our knowledge, the STZITD-GNN is the first work to use probabilistic framework for granular, multistep road-level crash forecasting. The use of Likelihood-based Loss minimization rather than the tradional regression ones enhances its uncertainty-aware predictive accuracy and reliability.
@@ 2 Literature Review
Predicting traffic crashes is essential for urban planning and mobility management. Initially, research in this area predominantly utilised traditional statistical methods and classical machine learning techniques, such as regression models (Chang,, 2005; Lord et al.,, 2005; Caliendo et al.,, 2007; Bergel-Hayat et al.,, 2013), Bayesian networks (Martin et al.,, 2009; Mujalli and De Oña,, 2011; Hossain and Muromachi,, 2012), tree-based models (Chong et al.,, 2004; Chang and Chen,, 2005; Wang et al.,, 2010; Lin et al.,, 2015), and the k-nearest-neighbour method (Sayed and Abdelwahab,, 1998; Tang and Gao,, 2005; Lv et al.,, 2009). These early studies were generally confined to small geographical areas and demonstrated limited effectiveness in capturing the nonlinear and dynamic spatio-temporal patterns of the contributing factors (Chang and Chen,, 2005; Anderson,, 2009; Zhang et al.,, 2014). In addition, these methods often analyse crash data in isolation, neglecting critical interdependencies between different locations (Wang et al.,, 2013; Wang et al., 2021a, ). As a result, their broader applicability to city-wide analysis with large data size was constrained, highlighting the need for more advanced models capable of addressing these limitations.
@@
Recent advances in deep learning techniques have shown promise in urban traffic crash prediction by incorporating spatio-temporal characteristics of the data. Convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph-based networks, such as graph-convolution networks (GCNs), have demonstrated their effectiveness in capturing intricate spatial patterns, capturing temporal dependencies, and modelling complex interactions (Wang and Cao,, 2021; Liu et al.,, 2020). As summarised in Table 1, Chen et al., (2016) were among the first to explore city level crash prediction employing a stack denoise autoencoder. They integrated human mobility GPS data and historical incident point data at the grid level to map the real-time crash situation in Tokyo. However, while their approach attempted to capture a larger spatial area, it did not consider urban geo-semantic information for precise and long-term crash prediction. Later, Chen et al., (2018) introduced a Stack Denoise Convolutional Autoencoder that incorporated spatial dependencies using stacked CNNs. Nonetheless, both of these studies overlooked the influence of temporal factors. Although RNNs are capable of handling temporal information, they are more suitable for short-term temporal learning (Ren et al.,, 2018). To better address the temporal aspect, Ren et al., (2018) used Long-Short-Term Memory (LSTM) to consider influential temporal factors between several locations. Their work represented the first attempt to incorporate temporal factors into deep learning models for the prediction of city-wide traffic crashes. Similarly, Moosavi et al., (2019) employed LSTM networks for country-wide crash prediction, demonstrating their efficacy in predicting a binary outcome. Building on the LSTM framework, Yuan et al., (2018) proposed a more advanced model known as the Convolutional Long Short-Term Memory (Hetero-ConvLSTM) neural network based on spatial heterogeneous data. This model aimed to jointly address spatial heterogeneity and capture temporal features in state-wide crash prediction. However, their approach relied on pre-defined moving windows based on the spatial features of Iowa, United States, limiting its generalisability to other cities. Bao et al., (2019) integrated CNNs, LSTM as well as ConvLSTM into a spatiotemporal convolutional long short-term memory network (STCL-Net). Their work presented a more practical and feasible approach for the prediction of city-wide traffic crashes. By combining the strengths of these models, STCL-Net effectively captured spatiotemporal dependencies in crash data and influential factors, leading to improved prediction performance.
@@
To capture local and global dynamics with hierarchical spatial information, Zhu et al., (2019) proposed the Deep Spatio-temporal Attention Learning Framework, which incorporates a spatio-temporal attention mechanism to effectively capture the dynamic impact of traffic crashes at different spatial levels. One notable aspect of their work was the emphasis on using real traffic administrative areas instead of manually divided grids. However, this approach does not incorporate spatial relationships, which limits its capacity to model the spatial heterogeneities that characterise traffic crashes. In a similar vein, Huang et al., (2019) introduced the Deep Dynamic Fusion Network (DFN) framework, which aggregates heterogeneous external factors across both spatial and temporal dimensions. Using a temporal aggregation layer, the DFN framework is designed to automatically capture external influences from latent temporal dimensions.
@@
Although the mentioned studies effectively captured temporal and spatial patterns in traffic crash data, yielding promising prediction outcomes, their lack of consideration of latent spatial correlations within urban structures can disrupt regional patterns in practical urban science and lead to prediction inaccuracies. Therefore, Graph-based networks have become a key method for city-wide traffic crash prediction by leveraging spatial graph features from actual urban networks (Xue et al.,, 2022; Cheng et al.,, 2022; Li et al., 2022a, ; Li et al., 2022b, ).
@@
Zhou et al., 2020b developed the RiskOracle framework by constructing an urban graph of regions and integrating the Differential Time-varying Graph Neural Network (DTGN) to identify dynamic correlations between historical traffic conditions and risks, enabling simultaneous predictions of traffic flow and crashes. To tackle zero-inflation in crash data, they introduced a data enhancement strategy based on a priori knowledge (PKDE), defining subregional crash labels for training. This approach adjusts zero crash values to a range between 0 and 1, maintaining the hierarchy of actual crash levels. Expanding on this foundation, Zhou et al., 2022b refined the DTGN framework by embedding an LSTM-based hierarchical sequence learning architecture, accommodating multiscale historical spatio-temporal data for nuanced long-term and short-term predictions. Wang et al., 2021a and Wang et al., 2021b both focus on the integration of geographical and semantic relationships through different graph structures. While Wang et al., 2021a used crash, road, and POI graphs to capture regional semantics and introduced modules for geographical and semantic spatio-temporal correlations, Wang et al., 2021b ) combined attention-based LSTM, CNN, and GCN for multi-view predictions of traffic crashes at various scales. Although both approaches tackled the zero-inflation issue with promising results, they share limitations in addressing dynamic geographical semantics and alignment issues in graph structures. In contrast, Trirat et al., (2023) proposed a multi-view graph neural network incorporating dynamic and static similarity information, offering a more dynamic approach to understanding traffic crashes. Using the Huber loss function, a parameter-based regression loss, their model aimed at robust regression that effectively accommodates zero inflation, showcasing an advanced strategy for precise prediction of urban traffic crashes.
@@
Recent attempts to apply graph-based deep learning for road-level crash prediction have shown promise, but also highlight areas for improvement. Yu et al., (2021) developed a Deep Spatiotemporal Graph Convolutional Network with a unique three-layer network, treating the road graph, spatio-temporal data, and embeddings separately. They tackled zero inflation through undersampling, balancing the dataset between risky and non-risky roads, and framing crash prediction as a binary classification. While this method effectively predicts crashes at the road level, the binary outcome overlooks the varying degrees of crash risk across roads. Similarly, Wu et al., (2022) designed a fully connected dynamic network to predict road-level crashes by focussing on dynamic spatio-temporal correlations. To mitigate zero inflation, they introduced a cost-sensitive learning module that enhances positive sample classification accuracy by considering the imbalance between zero and non-zero crash values. However, this work only considered limited road segments in New York City with binary crash occurrence outputs, highlighting the need for further development to capture the full spectrum of crash severity and provide a more detailed understanding of road-level crash scenarios.
@@
In summary, despite significant progress in urban traffic crash prediction yielding promising results, existing studies have yet to fully address uncertainty quantification in their predictions and provide multi-step crash assessments at the detailed road level.
@@
Table 1: Summary of Deep Learning Methods in Traffic Crash Prediction
@@
Chen et al., (2016) Stack Denoise Autoencoder Regions N/A Single Crash Risk Score
Ren et al., (2018) Long Short-term Memory network Regions N/A Multiple Crash Risk Score
Chen et al., (2018) Stack Denoise Convolutional Autoencoder Regions N/A Single Crash Risk Score
Yuan et al., (2018) Convolutional Long Short-Term Memory Regions N/A Multiple Crash Risk Score
Moosavi et al., (2019) Long Short-term Memory Network Regions N/A Single Crash Occurrence
Zhu et al., (2019) Deep Spatial-Temporal Attention Learning Framework Regions N/A Multiple Crash Risk Score
Huang et al., (2019) Dynamic Fusion Network Regions N/A Single Crash Occurrence
Bao et al., (2019) Spatiotemporal Convolutional LSTM Regions N/A Single Crash Risk Score
Zhou et al., 2020b DTGN with PKDE Regions Yes Single Crash Risk Score
Wang et al., 2021a GSNet with Weighted Loss Regions Yes Single Crash Risk Score
Wang et al., 2021b MVMT-ST Networks Regions Yes Single Crash Risk Score
Yu et al., (2021) Deep ST-GCN with Undersampling Roads Yes Single Crash Occurrence
Wu et al., (2022) MADGCN with Cost-sensitive Loss Roads Yes Multiple Crash Occurrence
Zhou et al., 2022b Enhanced DTGN Regions Yes Multiple Crash Risk Score
Trirat et al., (2023) MV-GNNs with Huber Loss Regions Yes Multiple Crash Risk Score